x^2 + x sqrt(x^2 + x) - 2 = 0
Find all real solutions. Please help as once again this one is getting the better of me!

Answer by edjones(8007)

x^2 + x sqrt(x^2 + x) - 2 = 0
x^2-2=-xsqrt(x^2+x)
x^4-4x^2+4=x^4+x^2 square each side.
5x^2-4=0
5x^2=4
x^2=4/5
x=+-2/sqrt(5)
x=+-2sqrt(5)/5
.
Ed